Please zoom in if you cannot see.

Part A A closely wound search coil has an area of 3.26 c2145 turns, and a resistance of 61 9 Ω It is connected to charge-measuring instrument whose resistance is 46 32 When the coil is rotated quicky from a position parallcl to a unitorm magnctic tield to one perpcnd cular to the teld, the instrurment indicales a chage of 3.46x10C What is the magnitude of the feld? B- Submit Provide Feedback
Answer #1

Given,

= Current I   =   e / R

= e   =   induced emf

= R =   total resistance   =   61.9 + 46.3 =   108.2 Ω

Area   A   =   3.26cm^2 =   3.26 * 10^-4 m^2

Then,

= dQ / dt   =   ( - dφ/dt) / R

= dQ   =   - dφ / R

=  - N * A * ΔB / R

Then,

Let the initial magnetic field   =   B

The final magnetic field   =   0

So,

= dQ   =   3.46 * 10^-5 =    - 145 * 3.26 * 10^-4 * (0 - B) / 108.2

= 3.46 * 10^-5 = 4.36 * 10^-4 * B

= B   =   3.46 * 10^-5 / 4.36 * 10^-4

=   0.0793 T